OpenStudy (bayan_610):

Which equation is equivalent to y=1/3x-5? A. Y-1=-3(x+2) B. Y+1=1/3(x+2) C. Y+1=1/3(x-12)

OpenStudy (mathmale):

If, by 1/3x, you mean (1/3)x, please use parentheses around the fraction 1/3 so as to leave no doubt about your intent. I am assuming that by "y=1/3x-5" you mean y=(x/3) - 5 or (1/3)x - 5. consider this as you attempt to choose the correct equivalent.

OpenStudy (bayan_610):

No that's how the answer is written the way I wrote it. @mathmale